Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2014-09-14

Note: Technical issue with pointing was resolved by engineering staff during daytime 15th Sept. Problem was caused by time server in NTT reporting incorrect year. Interim fix consists of using time signal from 3.6m for NTT, while faulty time server is repaired. No problems with pointing on subsequent night.


Observers:  Morgan Fraser, Chris Ashall
TAT: Joe Anderson, Annalisa de Cia
Data reducers: Nic Walton, Nadia Blagorodnova

Conditons:
No clouds visible at evening twilight, thin haze towards horizon - CLR
UT 23.00:    5 m/s wind, 35% humidity, pressure high
UT 03.00     Seeing measured at 1.7", pointing restrictions due to wind
UT 05:00     Closed due to wind until 07:00
UT 08:00     Thin clouds visible by eye in moonlight. THN

22.40    Twilight
             UVi skyflats

23.53    Attempted to acquire EG274, EG131, ASASSN-14gs, SN 2009ip. None of the fields were recognisable, and PESSTO_EFOSC_PHOT gives invalid WCS solutions (as checked against USNO) for all. SN 2009ip in particular is particularly distinctive, so do not believe this is observer error. After discussion with TO, we think there is a problem with the telescope pointing.

01.55    Problem continues, TO still attempting to fix.

02.48    Problem persists. Have measured the offset (~15') by solving for WCS of acquisition images, and comparing to expected RA and Dec. Now manually correcting all target coordinates for offset. OBs prefixed with "x" have an offset applied.

             EG274 std

03:05    ASASSN-14gm    Gr16

             Could not get a guide star - TO cannot apply measured offset to guide probe, and guide star catalog is showing incorrect position. Attempting to take short spectra with just tracking. Wind also increasing, getting close to pointing limit.

03:56    Could not take Gr#11 spectrum for ASASSN-14gm due to wind restrictions.
            
             LSQ14few      Classification, normal Ia around max light at z=0.06

05:09    Continuing problems with acquisition due to pointing errors. PESSTOEFOSCPHOT not working due to size of offsets.

05:10    CLOSED DUE TO WIND

07:03    REOPENED
      
             LSQ14fep    Classification, normal Ia, between a few days pre-max and max, at z=0.06

08:00   ASASSN-14ha      Followup, Gr#11 and Gr#16.

08:45    Telesecope error, rebooting

08:57    
Weak features seen in ASASSN-14ha Gr#11 spectrum (between 3500 and 4500 Ang), so taking second Gr#11 exposure to increase S/N here.


Comments